### v3

# RapVerse: Coherent Vocals and Whole-Body Motion Generation from Text

Abstract
In this work, we introduce a challenging task for simultaneously generating 3D holistic body motions and singing vocals directly from textual lyrics inputs, advancing beyond existing works that typically address these two modalities in isolation. To facilitate this, we first collect the RapVerse dataset, a large dataset containing synchronous rapping vocals, lyrics, and high-quality 3D holistic body meshes. With the RapVerse dataset, we investigate the extent to which scaling autoregressive multimodal transformers across language, audio, and motion can enhance the coherent and realistic generation of vocals and whole-body human motions. For modality unification, a vector-quantized variational autoencoder is employed to encode whole-body motion sequences into discrete motion tokens, while a vocal-to-unit model is leveraged to obtain quantized audio tokens preserving content, prosodic information and singer identity. By jointly performing transformer modeling on these three modalities in a unified way, our framework ensures a seamless and realistic blend of vocals and human motions. Extensive experiments demonstrate that our unified generation framework not only produces coherent and realistic singing vocals alongside human motions directly from textual inputs, but also rivals the performance of specialized single-modality generation systems, establishing new benchmarks for joint vocal-motion generation. Video demonstration and more information can be found on the project page [^1] https://jiabenchen.github.io/RapVerse/..
Figure 1: RapVerse. We present a unified text-conditioned multi-modality generation framework, for jointly generating holistic body motions and singing vocals from textual lyrics inputs only. Note that the corresponding video frames are just shown for reference.
1 Introduction
In the evolving landscape of multi-modal content generation in terms of sound and motion, significant strides have been made in individual modalities, including text-to-music [schneider2023mo, agostinelli2023musiclm, huang2023noise2music], text-to-vocal [liu2022diffsinger], text-to-motion [guo2022generating, zhang2023t2m, chen2023executing, jiang2024motiongpt, lu2023humantomato], and audio-to-motion [yi2023generating, habibie2021learning, liu2023emage] generation. These developments have paved the way for creating more dynamic and interactive digital content. Despite these advancements, existing works predominantly operate in silos, addressing each modality in isolation. However, there’s strong psychological evidence that for human beings, the generation of sound and motion are highly related and coupled [lakoff1980metaphors]. A unified system for joint generation allows for a more expressive and nuanced communication of emotions, intentions, and context, where the generation of one modality could guide and assist the other in a coherent and efficient way.
In this paper, we tackle a crucial problem: can a machine not only sing with emotional depth but also perform with human-like expressions and motions? We propose a novel task for generating coherent singing vocals and whole-body human motions (including body motions, hand gestures, and facial expressions) simultaneously, see Fig. 1. This endeavor holds practical significance in fostering more immersive and naturalistic digital interactions, thereby elevating virtual performances, interactive gaming, and the realism of virtual avatars.
An important question naturally arises: what constitutes a good model for unified generation of sound and motion? Firstly, we consider textual lyrics as the proper form of inputs for the unified system, since text provides a highly expressive, interpretable flexible means of conveying information by human beings, and could serve as a bridge between various modalities. Previous efforts explore scores [liu2022diffsinger], action commands [zhang2023t2m, chen2023executing, jiang2024motiongpt], or audio signals [yi2023generating] as inputs, which are inferior to textual inputs in terms of semantic richness, expressiveness and flexible integration of different modalities.
Secondly, we reckon that a joint generation system that could produce multi-modal outputs simultaneously is better than a cascaded system that executes the single-modal generation sequentially. A cascaded system, combining a text-to-vocal module with a vocal-to-motion module, risks accumulating errors across each stage of generation. For instance, a misinterpretation in the text-to-vocal phase can lead to inaccurate motion generation, thereby diluting the intended coherence of the output. Furthermore, cascaded architectures necessitate multiple training and inference phases across different models, substantially increasing computational demands.
To build such a joint generation system, the primary challenges include: 1) the scarcity of datasets that provide lyrics, vocals, and 3D whole-body motion annotations simultaneously; and 2) the need for a unified architecture capable of coherently synthesizing vocals and motions from text. In response to these challenges, we have curated RapVerse, a large-scale dataset featuring a comprehensive collection of lyrics, singing vocals, and 3D whole-body motions. Despite the existence of datasets available for text-to-vocal [liu2022diffsinger, huang2021multi, duan2013nus, sharma2021nhss], text-to-motion [plappert2016kit, mahmood2019amass, guo2022generating, lin2024motion], and audio-to-motion [cao2017realtime, habibie2021learning, ginosar2019learning, fanelli20103, cudeiro2019capture, wuu2022multiface], the landscape lacks a unified dataset that encapsulates singing vocals, whole-body motion, and lyrics simultaneously. Most notably, large text-to-vocal datasets [huang2021multi, zhang2022m4singer] are predominantly in Chinese, limiting their applicability for English language research and lacking any motion data. And text-to-motion datasets [plappert2016kit, guo2022generating, lin2024motion] typically focus on text descriptions of specific actions paired with corresponding motions without audio data, often not covering whole body movements. Moreover, audio-to-motion datasets [liu2022diffsinger, lu2023co] focus primarily on speech rather than singing. A comparison of existing related datasets is demonstrated in Table. 1. The RapVerse dataset is divided into two distinctive parts to cater to a broad range of research needs: 1) a Rap-Vocal subset containing a large number of pairs of vocals and lyrics, and 2) a Rap-Motion subset encompassing vocals, lyrics, and human motions. The Rap-Vocal subset contains 108.44 hours of high-quality English singing voice in the rap genre without background music. Paired lyrics and vocals are crawled from the Internet from 32 singers, with careful cleaning and post-processing. On the other hand, the Rap-Motion subset contains 26.8 hours of rap performance videos with 3D holistic body mesh annotations in SMPL-X parameters [pavlakos2019expressive] using the annotation pipeline of Motion-X [lin2024motion], synchronous singing vocals and corresponding lyrics.
With the RapVerse dataset, we explore how far we can push by simply scaling autoregressive multimodal transformers with language, audio, and motion for a coherent and realistic generation of vocals and whole-body human motions. To this end, we unify different modalities as token representations. Specifically, three VQVAEs [van2017neural] are utilized to compress whole-body motion sequences into three-level discrete tokens for head, body, and hand, respectively. For vocal generation, previous works [min2021meta, donahue2020end, liu2022diffsinger, min2021meta] share a common paradigm, producing mel-spectrograms of audio signals from input textual features and additional music score information, following with a vocoder [oord2016wavenet, valin2019lpcnet, yang2021multi] to reconstruct the phase. We draw inspiration from the speech resynthesis domain [polyak2021speech], and learn a self-supervised discrete representation to quantize raw audio signal into discrete tokens while preserving the vocal content and prosodic information. Then, with all the inputs in discrete representations, we leverage a transformer to predict the discrete codes of audio and motion in an autoregressive fashion. Extensive experiments demonstrate that this straightforward unified generation framework not only produces realistic singing vocals alongside human motions directly from textual inputs but also rivals the performance of specialized single-modality generation systems.
To sum up, this paper makes the following contributions:
  • We collect RapVerse, a large dataset featuring synchronous singing vocals, lyrics, and high-quality 3D holistic SMPL-X parameters.
  • We design a simple but effective unified framework for the joint generation of singing vocals and human motions from text with a multi-modal transformer in an autoregressive fashion.
  • To unify representations of different modalities, we employ a vocal-to-unit model to obtain quantized audio tokens and utilize compositional VQVAEs to get discrete motion tokens.
  • Experimental results show that our framework rivals the performance of specialized single-modality generation systems, setting new benchmarks for joint generation of vocals and motion.
2 Related Work
2.1 Text to Vocal Generation
Text-to-audio Dataset. There exists several singing vocal datasets, yet they each has constraints. For instance, PopCS [liu2022diffsinger] and OpenSinger [huang2021multi] are limited to Chinese, while NUS-48E [duan2013nus] and NHSS [sharma2021nhss] have only a few hours of songs. Nevertheless, JVS-MuSiC [tamaru2020jvs] and NUS-48E [duan2013nus] offer a few hours of songs from dozens of singers, whereas OpenSinger [huang2021multi] provides a more extensive collection with tens of hours from a single singer. Notably, our dataset represents the first to be specifically curated for rap songs of multiple singers with 108 hours.
Text-to-vocal Models. Recent advancements in text-to-speech (TTS) models, including WavNet [oord2016wavenet], FastSpeech 1 and 2 [ren2019fastspeech, ren2020fastspeech], and EATS [donahue2020end], have significantly improved the quality of synthesized speech. However, singing voice synthesis (SVS) presents a greater challenge due to its reliance on additional musical scores and lyrics. Recent generation models [liu2022diffsinger, he2023rmssinger, he2023rmssinger, zhang2022wesinger, kim2023muse] perform excellently in generating singing voices. However, their performance tends to deteriorate when encountering out-of-distribution data. To handle this, StyleSinger [zhang2023stylesinger] introduces a Residual Style Adaptor and an Uncertainty Modeling Layer Normalization to handle this.
Dataset Language #Singers Hours
NUS-48E [duan2013nus] English 12 1.91
NHSS [sharma2021nhss] English 10 7
JVS-MuSiC [tamaru2020jvs] Japanese 100 2.28
Tohoku Kiritan [ogawa2021tohoku] Japanese 1 1
PopCS [liu2022diffsinger] Chinese 1 5.89
OpenSinger [huang2021multi] Chinese 66 50
Opencpop [wang2022opencpop] Chinese 1 5.25
M4Singer [zhang2022m4singer] Chinese 20 29.77
Ours English 32 108.44
Dataset Modality Whole Hours
KIT-ML [plappert2016kit] Text
11.2
AMASS [mahmood2019amass] Text
40.0
HumanML3D [guo2022generating] Text
28.6
Motion-X [lin2024motion] Text
127.1
HA2G [liu2022diffsinger] Audio
33
Yoon et.al [lu2023co] Audio
30
Speech2Gesture [ginosar2019learning] Audio
144
Talkshow [yi2023generating] Audio
27.0
Ours Text-Audio
26.8
Table 1: Comparison of Audio and Motion Datasets. (a) compares our Rap-Vocal Subset with existing singing vocal datasets. (b) compares our Rap-Motion Subset with existing motion datasets. : “Whole” means whole-body motion including face, body, and hand.
Figure 2: RapVerse dataset collection pipeline. There are two pathways for rap-vocal and rap-motion subsets, respectively.
2.2 Text to Motion Generation
Text-to-motion Dataset. Current text-motion datasets, such as KIT [plappert2016kit], AMASS [mahmood2019amass], and HumanML3D [guo2022generating], are constrained by limited data coverage, typically spanning only tens of hours and lacking whole-body motion representation. To overcome these shortcomings, Motion-X [lin2024motion] was developed, providing an extensive dataset that encompasses whole-body motions.
Text-to-motion Models. Text-to-motion models [petrovich2022temos, zhang2024motiondiffuse, guo2022tm2t, guo2022generating, lu2023humantomato, chen2023executing] have gained popularity for their user convenience. Recent advancements have focused on diffusion models [zhang2024motiondiffuse, kim2023flame, tevet2022human], which, unlike deterministic generation models [ahuja2019language2pose, ghosh2021synthesis], enable finer-grained and diverse output generation. Chen et al. [chen2023executing] proposed a motion latent-based diffusion model to enhance generative quality and reduce computational costs. However, the misalignment between natural language and human motions presents challenges. Lu et al. [lu2023humantomato] introduced the Text-aligned Whole-body Motion generation framework to address these issues by generating high-quality, diverse, and coherent facial expressions, hand gestures, and body motions simultaneously.
2.3 Audio to Motion Generation
Audio-to-motion Dataset. Co-Speech Datasets can be classified into 1: pseudo-labeled (PGT) and 2: motion-captured (mocap). PGT datasets [cao2017realtime, habibie2021learning, ginosar2019learning, yi2023generating] utilize disconnected 2D or 3D key points to represent the body, allowing extracting at a lower cost but with limited accuracy. On the other hand, mocap datasets provide annotations for limited body parts, with some focusing solely on the head [fanelli20103, cudeiro2019capture, wuu2022multiface] or the body [takeuchi2017creating, ferstl2018investigating]. Differing from them, BEATX [liu2023emage] contains mesh data of both head and body. However, these datasets typically focus on speech to motion generation. Our RapVerse, in contrast, contains paired text-vocal-motion data, enabling simultaneous motion and vocal generation.
Audio-to-motion Models. Audio-to-motion models [taylor2017deep, shen2023difftalk, zhang2023sadtalker, liu2023emage, yi2023generating, alexanderson2023listen, bian2024adding, chen2024diffsheg] aim to produce human motion from audio innputs. Recognizing the intricate relationship between audio and human face, TalkSHOW [yi2023generating] separately generates face parts and other body parts. However, this approach exhibits limitations, such as the absence of generating lower body motion. Motioncraft [bian2024adding] proposes a DiT structure for motion generation. EMAGE [liu2023emage] utilizes masked gestures together with vocals for a generation. Our model, however, goes a step further by being the first to generate paired audio-motion data directly from text.
3 RapVerse Dataset
In this section, we introduce RapVerse, a large rap music motion dataset containing synchronized singing vocals, textual lyrics and whole-body human motions. A comparison of our dataset with existing datasets is shown at Table. 1. The RapVerse dataset is divided into two subsets to cater to a broad range of research needs: a Rap-Vocal subset and a Rap-Motion subset. The overall collection pipeline of RapVerse is shown at Fig. 2.
3.1 Rap-Vocal Subset
The Rap-Vocal subset contains 108.44 hours of high-quality English singing voice in the rap genre with paired lyrics. We will introduce each step in detail.
Data Crawling. In a bid to obtain a large number of rap songs and corresponding lyrics from the Internet, we utilize Spotdl and Spotipy to collect songs, lyrics, and metadata of different rap singers. To ensure the quality of the dataset, we perform cleaning on the crawled songs by removing songs with misaligned lyrics and filtering out songs that are too long or too short.
Vocal and Background Music Separation. Since the crawled songs are mixed with rapping vocals and background music, and we aim to synthesize singing vocals from separated clean data, we utilize Spleeter [hennequin2020spleeter], the state-of-the-art open-source vocal-music source separation tool to separate and extract rap vocal voices and accompanying background music from the collected songs. Following [ren2020deepsinger], we normalize the loudness of the vocal voices to a fixed loudness level.
Vocal Data Processing. The raw crawled lyrics from the Internet are in inconsistent formats, we conduct data cleaning on the lyrics by removing meta information (singer, composer, song name, bridging words, and special symbols). To ensure that the lyrics are aligned with the singing vocals, we collect lyrics only with the correct timestamps of each sentence, and we separate each song into around 10-second to 20-second segments for model training.
3.2 Rap-Motion Subset
The Rap-Motion subset contains 26.8 hours of rap performance videos with 3D holistic body mesh annotations in SMPL-X parameters [pavlakos2019expressive], synchronous singing vocals, and corresponding lyrics. We introduce the collection pipeline of this subset as follows.
Data Crawling. We crawled over 1000 studio performance videos from YouTube under the Common Creative License. We filter out low-quality videos manually to ensure the videos meet the following criteria: stable camera work, performers centered in the frame, clear visibility of the performer’s entire body to capture detailed motion data, and high-quality audio for accurate vocal analysis.
Audio Data Processing. Similar to the Rap-Vocal subset, we leverage Spleeter [spleeter2020] to isolate singing vocals from accompanying music. Given that YouTube videos typically lack paired lyrics, we employ an ASR model, Whisper [radford2023robust], to accurately transcribe vocals into corresponding text.
Video Data Processing. To ensure the collection of high-quality video clips for motion annotation, we implemented a semi-automatic process to filter out undesirable content, such as advertisements, transition frames, changes in shots, and flashing lights. Initially, we applied YOLO [redmon2018yolov3] for human detection to discard frames where no humans were detected. Subsequently, we utilized RAFT [teed2020raft] to assess the motion magnitude, employing a threshold to eliminate frames affected by camera instability. We then perform meticulous manual curation on the extracted clips, retaining only those of the highest quality. Finally, we follow the pipeline of the optimized-based method Motion-X [lin2024motion] to extract 3D whole-body meshes from monocular videos. Specifically, we adopt the SMPL-X [pavlakos2019expressive] for motion representations, given a T-frame video clip, the corresponding pose states $\mathcal{M}$ are represented as:
$\mathcal{M}=\{\mathcal{M}_{f},\mathcal{M}_{b},\mathcal{M}_{h},\zeta,\epsilon\}$ (1)
where $\mathcal{M}_{f}\in\mathbb{R}^{T\times 3},\mathcal{M}_{b}\in\mathbb{R}^{T\times 63}$ and $\mathcal{M}_{h}\in\mathbb{R}^{T\times 90}$ denotes jaw poses, body poses and hand poses, respectively. $\zeta\in\mathbb{R}^{T\times 100}$ and $\epsilon\in\mathbb{R}^{T\times 3}$ are the facial expression and global translation.
Figure 3: Pipeline overview. We first pre-train all tokenizers on vocal-only and motion-only data. After we have pretrained the modality tokenizers, we can unify text, vocal, and motion in the same token space. We adopt a mixing organizing algorithm for input tokens to align via the temporal axis. These mixed input tokens are fed into the large Text-Motion-Audio foundation model to train on token prediction tasks, guided by the encoded features from textual input.
4 Method
Given a piece of lyrics text, our goal is to generate rap-style vocals and whole-body motions, including body movements, hand gestures, and facial expressions that resonate with the lyrics. With the help of our RapVerse dataset, we propose a novel framework that not only represents texts, vocals, and motions as unified token forms but also integrates token modeling in a unified model. As illustrated in Fig. 3, our model consists of multiple tokenizers for motion (Sec. 4.2) and vocal (Sec. 4.3) token conversions, as well as a general Large Text-Motion-Audio Foundation Model (Sec. 4.4) that targets for audio token synthesize and motion token creation, based on rap lyrics.
4.1 Problem Formulation
Let $\mathcal{L}$ be the rap lyric, such as “I going to make it a flashback set a picture from way back.”, our model will compose the text-related vocal $\mathcal{V}\in\mathbb{R}^{t\times d_{v}}$ and whole-body motion $\mathcal{M}\in\mathbb{R}^{t\times d_{m}}$, where $t$, $d_{v}$, $d_{m}$ denote the temporal length and the feature dimensions for vocal and motion in every time unit, respectively. Modality tokenizer encoders $\phi^{\mathcal{X}}(\mathcal{T}\mid\mathcal{X})$ and decoders $\varphi^{\mathcal{X}}(\mathcal{X}\mid\mathcal{T})$ will bridge each modality with the token format $\mathcal{T}$ utilizing different tokenizers, where $\mathcal{X}\in\{\mathcal{L},\mathcal{V},\mathcal{M}\}$. The large foundation model $\psi^{\mathcal{X}}$ will incorporate tokens in unified modeling. The whole problem of coherent vocal and model generation via lyric text can be formulated as follows:
$\Theta^{*}=\mathop{\arg\max}_{\Theta}\mathbb{P}_{\Theta}(\mathcal{V},\mathcal{M}\mid\mathcal{L}),$ (2)
where $\Theta$ denotes the whole model parameters and $\mathbb{P}_{\Theta}(\cdot)$ denotes the model distributions.
4.2 Motion VQ-VAE Tokenizer
Similar to prior arts, a motion VQ-VAE is adopted to convert 3D human motions into discrete tokens, where the tokenizer encoder $\phi$ generates motion tokens enriched with high-frequency information, while the decoder $\varphi$ seamlessly reconstructs discrete tokens into continuous motion sequences. Specifically, the tokenizer will maintain a learnable codebook $\mathcal{C}$, which firstly encodes the input motion features $\mathcal{M}$ to vectors $z^{\mathcal{M}}$, and then the motion quantizer $\mathcal{Q}^{\mathcal{M}}$ will look up the nearest neighbor in codebook $\mathcal{C}^{\mathcal{M}}=\{c_{k}^{\mathcal{M}}\}_{k=1}^{d_{c}}$ for the motion representation. The quantized code can be calculated as
$Q(z^{\mathcal{M}};\mathcal{C}^{\mathcal{M}})=\mathop{\arg\min}_{k}\|z^{\mathcal{M}}-c_{k}^{\mathcal{M}}\|_{2}.$ (3)
The tokenizer decoder $\varphi$ will project the quantized code back into the human motion after quantization. The reconstruction objective in the tokenizer training stage can be represented as
$\mathbb{E}_{\mathcal{T}\sim\phi_{\mathcal{M}}(\mathcal{T}\mid\mathcal{M})}\left[\log\varphi_{\mathcal{M}}(\mathcal{M}\mid\mathcal{T})\right].$ (4)
As the whole-body motion $\mathcal{M}$ can be decomposed into face $\mathcal{M}_{f}$, body $\mathcal{M}_{b}$, and hand $\mathcal{M}_{h}$, where $\mathcal{M}\in\mathbb{R}^{t\times d},d_{f}+d_{b}+d_{h}=d_{m}$, we build three tokenizers to obtain tokens for these three parts.
4.3 Vocal2unit Audio Tokenizer
Overall, we leverage the self-supervised framework [polyak2021speech] in speech resynthesis domain to learn vocal representations from the audio sequences. Specifically, we train a Vocal2unit audio tokenizer to build a discrete tokenized representation for the human singing voice. The vocal tokenizer consists of three encoders and a vocoder. The encoders include three different parts: (1) the semantic encoder; (2) the F0 encoder; and (3) the singer encoder. We will introduce each component of the model separately.
Semantic Encoder. In order to extract the semantic information of the audio into discrete tokens, we use a pre-trained Hubert encoder [hsu2021hubert] as the main component of our semantic encoder network $E^{h}$. Given a raw audio sequence $v$, the output of the Hubert encoder is a continuous representation. Then we apply a K-means algorithm on the representations to finally generate discrete units $z^{\mathcal{S}}=\{z^{\mathcal{S}}_{i}\}_{i=1}^{L_{s}}$ where $z^{\mathcal{S}}_{i}\in\{1,2,\cdots,K\}$. $L_{s}$ is the length of hubert token sequence, and $K$ represents the number of clusters in K-means algorithm.
F0 Encoder. We use the YAAPT algorithm [kasi2002yet] to extract the F0 from the input audio signal $v$, which is then fed into an encoder $E_{p}$ to generated low-frequency F0 pitch representation. The encoder $E^{p}$ is trained using a VQ-VAE framework with Exponential Moving Average updates similar to [dhariwal2020jukebox], maintaining a learnable codebook $\mathcal{C}^{p}=\{c_{k}^{p}\}_{k=1}^{d_{c}}$, where $d_{c}$ is the number of codes in the codebook. The raw F0 sequence is first preprocessed according to the singer, converted into a sequence of latent vectors, then mapped to its nearest neighbors in the codebook $c^{p}$, which could be represented using integer number $k\in\{1,2,\cdots,d_{c}\}$ as its index in the codebook. The index sequence will be further formatted as the token ¡pitch_k¿ to produce pitch token sequence $z^{\mathcal{P}}=\{z^{\mathcal{P}}_{i}\}_{i=1}^{L_{p}}$, where $L_{p}$ is the length of the pitch token sequence. Finally, the decoder will reconstruct the F0 representation leveraging the pitch token sequence.
Singer Encoder. As the last component of our encoders, the singer encoder $E_{si}$ from [heigold2016end, polyak2021speech] is used to extract the Mel-spectrogram from the raw audio sequence and to output a singer representation $z^{\mathcal{I}}\in\mathbb{R}^{256}$. The singer embedding only depends on the singer and is trained globally.
Vocoder. In order to decode the vocal signal from the discrete tokens, we adopt a modified version of the HiFi-GAN neural vocoder similar to the one in [polyak2021speech], consisting of a generator and multiple discriminators. The generator takes the encoded discrete representation $z^{\mathcal{S}}$ and $z^{\mathcal{P}}$, and the singer embedding vector $z^{\mathcal{I}}$ as inputs. After a lookup table and a series of blocks composed of transposed convolution and a residual block with dilated layers, it finally outputs waveform audio. The discriminators include multiple sub-discriminators, analyzing on different scales and different periodic structures.
4.4 General Auto-regressive Modeling
Model Architectures. The whole model of our general modeling pipeline consists of several pre-trained vocal and motion tokenizers and a large Text-Motion-Audio foundation model. The vocal and motion tokenizers are based on VQ-VAE architectures, enabling us to represent these two modalities as discrete tokens. We also adopt a T5-Tokenizer [raffel2020exploring] to convert lyric text into tokens, so that all three modalities $\mathcal{X}\in\{\mathcal{L},\mathcal{V},\mathcal{M}\}$ are united in the token space $\mathcal{T}^{\mathcal{X}}=\{t_{k}^{\mathcal{X}}\}_{k=1}^{\mathcal{N}^{\mathcal{X}}}$, where $\mathcal{N}^{\mathcal{X}}$ is the token sequence length for each modality. The large foundation model, a temporal transformer based on the decoder-only architecture, performs the next-token-prediction task based on textual features encoded by the T5-Encoder.
Specifically, each token is generated after the start token by the probability distribution from the foundation model $p_{\delta}(t^{\mathcal{X}};h^{\mathcal{L}})=\prod\limits_{i}p_{\delta}(t_{i}^{\mathcal{X}}\mid t_{<i}^{\mathcal{X}};h_{\mathcal{L}})$. During the training period, we calculate the cross-entropy loss to measure the likelihood between ground truth tokens and predicted token probabilities as
$\sum\limits_{i=1}^{\mathcal{N^{\mathcal{X}}}}\log p_{\delta}(t^{\mathcal{X}}_{i};h^{\mathcal{L}}).$ (5)
After optimizing via this training objective, our model learns to predict the next token, which can be decoded into different modality features. This process is similar to text word generation in language models, while the “word” in our method such as ¡face_0123¿, does not have explicit semantic information, but can be decoded into continuous modality features.
Multimodality Tokens Organization. Since we have designed a coherent vocal and motion generation pipeline with multiple token types in our method, including hubert and pitch tokens in rap vocals, and face, body, hand in whole-body motions, the organization method of all these tokens in one sentence is important for training. Similar to previous works [lu2023humantomato, jiang2024motiongpt], we adopt an interleaved style within each modality. Specifically, with the input whole-body motion tokens $\mathcal{T}^{\mathcal{M}_{s}}=\{t^{\mathcal{M}_{s}}_{k}\}_{k=1}^{\mathcal{N}^{\mathcal{M}_{s}}}$, where $s=f,b,h$ which denotes face, body and hand respectively, and $\mathcal{N}^{\mathcal{M}_{s}}$ denotes the token length for each part. Therefore the motion token sequence can be formulated as $\mathcal{T}^{\mathcal{M}}=\{t^{\mathcal{M}_{f}}_{1},t^{\mathcal{M}_{b}}_{1},t^{\mathcal{M}_{h}}_{1},t^{\mathcal{M}_{f}}_{2},$...\}. Similarly, the vocal tokens consist of the hubert tokens $t^{\mathcal{V}_{h}}$ and pitch tokens $t^{\mathcal{V}_{p}}$, which can be formulated as $\mathcal{T}^{\mathcal{V}}=\{t^{\mathcal{V}_{h}}_{1},t^{\mathcal{V}_{p}}_{1},t^{\mathcal{V}_{h}}_{2},t^{\mathcal{V}_{p}}_{2},$...\}. This interleaved way of organization allows us to align the timesteps of each part. When performing coherent vocal and motion generation, we mix these two modality sequences sequentially with modality-wise start token as $\mathcal{T}=\{t_{start}^{V},\mathcal{T}^{V},t_{start}^{M},\mathcal{T}^{M}\}$. Note that we put vocal sequences before motion sequences because 1) the vocals are more related to the lyrics, and 2) motion tokens have a relationship with previously generated vocals, e.g. face motions with lips movement are directly related to vocals. Joint training of mixed modalities allows the transformer to capture the temporal dependencies and dynamics inherent in vocal and motion data.
Inference and Decoupling. In the inference stage, we use different start tokens to specify which modality to generate. The textual input is encoded as features to guide token inference. We adopt a top-k algorithm to control the diversity of the generated content by adjusting the temperature, as generating vocals and motions based on lyrics is a creation process with multiple possible answers. After token prediction, a decoupling algorithm is used to process output tokens to make sure tokens from different modalities are separated and temporally aligned. These discrete tokens will be further decoded into text-aligned vocals and motions.
5 Experiments
In this section, we evaluate our proposed model on our proposed benchmark designed for joint vocal and whole-body motion generation from textual inputs.
5.1 Experimental Setup
Metrics. To evaluate the generation quality of singing vocals, we utilize the Mean Opinion Score (MOS) to gauge the naturalness of the synthesized vocal. For motion synthesis, we evaluate the generation quality of the body hand gestures and the realism of the face, respectively. Specifically, for gesture generation, we use Frechet Inception Distance (FID) based on a feature extractor from [guo2022generating] to evaluate the distance of feature distributions between the generated and real motions, and Diversity (DIV) metric to assess the motions diversity. For face generation, we compare the vertex MSE [xing2023codetalker] and the vertex L1 difference LVD [yi2023generating]. Finally, we adopt Beat Constancy (BC) [li2021ai] to measure the synchrony of generated motion and singing vocals.
Baselines. We compare the vocal generation quality with the state-of-the-art vocal generation method DiffSinger [liu2022diffsinger]. And we also adapt the text-to-speech model FastSpeech2 [ren2020fastspeech] for vocal generation. For motion generation, we compare our method with both text-to-motion methods and audio-to-motion methods. For text-to-motion methods, since there is no existing open-sourced work for text to whole-body motion generation, we compare with transformer-based T2M-GPT [zhang2023t2m] and MLD [chen2023executing] for body generation. For the audio-to-motion generation, we compare with Habibie et al. [habibie2021learning], Talkshow [yi2023generating], Motioncraft [bian2024adding] and EMAGE [liu2023emage]. We report all the results on RapVerse with an 85%/7.5%/7.5% train/val/test split.
[…]
Text-to-Motion
T2M-GPT [zhang2023t2m] 23.45 11.75 - - -
MLD [chen2023executing] 26.34 12.15 - - -
Audio-to-Motion
Habibie et al. [habibie2021learning] 32.14 10.08 0.476 2.13 9.54
Talkshow [yi2023generating] 18.23 13.14 0.482 2.05 9.20
EMAGE [liu2023emage] 21.18 12.65 0.488 1.96 8.45
MotionCraft [bian2024adding] 17.75 13.76 0.482 2.06 9.23
Cascaded Result 23.42 12.87 0.479 2.09 9.38
Text-to-Audio+Motion
Ours 17.58 14.08 0.485 2.03 7.23
[…]
GT 4.45 ± 0.06
Reconstruction 4.02 ± 0.08
FastSpeech2 [ren2020fastspeech] 3.41 ± 0.18
DiffSinger [liu2022diffsinger] 3.72 ± 0.12
Ours 3.64 ± 0.15
Table 2: Quantitative results of generated motion and vocal. (a) compares with different motion generation baselines, best result is shown in bold and the second-best result is shown in underline. (b) compares with text-to-vocal generation baselines. : The MOS is calculated with 95% confidence intervals of song samples.
5.2 Main Results Analysis
Evaluations on joint vocal and whole-body motion generations. We compared both text-driven and audio-driven motion generation baselines in Table. 2 (a). To be noted, our setting is different from all existing methods in the following ways. First, we use rap lyrics as our textual input instead of motion textual descriptions, which contain direct action prompt words, such as walk and jump; Second we use text to jointly generate both audio and motion, instead of using audio to generate motion as audio-driven methods did. As is demonstrated, our model rivals with both text-to-motion and audio-to-motion methods in terms of metrics measuring body motion quality and face motion accuracy.
[…]
GT 0 14.53 0.499 0 0
Pretrained LLM 48.25 12.65 0.462 3.28 12.15
Single motion token 19.15 12.75 0.477 2.18 10.12
Ours 17.58 14.08 0.485 2.03 7.23
Table 3: Ablation study. We compare with common designs in motion generation frameworks.
Figure 4: User study on generated motion.
Furthermore, the cornerstone of our approach lies in the simultaneous generation of vocals and motion, aiming to achieve temporal alignment between them. This objective is substantiated by our competitive results on the BC metric, which assesses the synchrony between singing vocals and corresponding motions, underscoring our success in closely synchronizing the generation of these two modalities. For the cascaded system, we integrate the text-to-vocal model DiffSinger with the audio-to-motion model Talkshow. Compared with the cascaded system, our joint-generation pipeline demonstrates superior outcomes while also reducing computational demands during both training and inference phases. In the cascaded architectures, errors tend to accumulate through each stage. Specifically, if the text-to-vocal module produces unclear vocals, it subsequently hampers the audio-to-motion model’s ability to generate accurate facial expressions that align with the vocal content.
To better evaluate the generation quality, we conducted an additional user study (see Fig. 4) using ten 10-second-long videos from the RapVerse dataset, sampled from the test set. We recruited 20 participants to evaluate four aspects: the holistic realism of human motion, expression-vocal synchronism, gesture-vocal synchronism, and holistic diversity, using real data as a reference. Participants were asked to sort shuffled videos generated by different methods based on these criteria. As shown in the results, our model achieved the highest scores in human motion realism, gesture-vocal synchronism, and holistic diversity when compared with T2MGPT (text-to-motion) and Talkshow (audio-to-motion).
Evaluations on vocal generations. We compare our method against other state-of-the-art text-to-vocal generation baselines in Table. 2 (b). While our unified model is trained to simultaneously generate vocals and motion, a task considerably more complex than generating vocals alone, its vocal generation component still manages to achieve results comparable to systems designed solely for vocal generations.
5.3 Ablation Study
We present the outcomes of our ablation study in Table. 3. Initially, we explored the integration of a pre-trained large language model [raffel2020exploring] for multi-modality generation, akin to the approach in  [jiang2024motiongpt]. However, the efficacy of utilizing pre-trained language models significantly lags behind our tailored design, underscoring that pre-training primarily on linguistic tokens does not facilitate effective prediction across multiple modalities like vocal and motion. Additionally, we study the impact of our compositional VQ-VAEs on motion generation. In contrast, a baseline employing a single VQVAE for the joint quantization of facial, body, and hand movements was implemented. This approach led to a noticeable degradation in performance, particularly marked by a -2.89 decrease in LVD. This decline can be attributed to the preponderance of facial movements in a singer’s performance. Using a single VQ-VAE model for full-body dynamics compromises the detailed representation of facial expressions, which are crucial for realistic and coherent motion synthesis.
6 Conclusion
In this work, we present a new framework for the simultaneous generation of 3D whole-body motions and singing vocals, directly from textual lyrics. To address this challenging task, we first collect RapVerse, a large dataset containing synchronous rap vocals, alongside lyrics and 3D whole-body motions. Utilizing RapVerse, we demonstrate that simply scaling autoregressive transformers across language, audio, and motion yields a coherent generation of singing vocals and 3D holistic human motions. We anticipate that this work will inspire novel avenues in the joint modeling of text, audio, and motion.